OpenStudy (anonymous):

show that integration of cosec 2x dx limit 1/6π to 1/3π = 1/2ln3

OpenStudy (anonymous):

@tkhunny @thomaster @SithsAndGiggles @Compassionate guys anyone help me please?

OpenStudy (tkhunny):

This? \(\int\limits_{\pi/6}^{\pi/3}\;\csc^{2}(x)\;dx\;=\;\dfrac{1}{2}\ln(3)\)?? Have you considered a substitution, maybe \(u = \cot(x)\)?

OpenStudy (anonymous):

please help me with this I'm dying trying to solve this :(

OpenStudy (anonymous):

@phi @mathmath333 @aum @kohai help me anyone :((((((

OpenStudy (phi):

I would re-write it using sin. Can you do that ?

OpenStudy (anonymous):

\[1/\sin2x\]

OpenStudy (phi):

now we need an idea.

OpenStudy (anonymous):

I used the double angle formula and I got 1/2sinxcosx but how to proceed from there?

OpenStudy (phi):

yes, I looked at that. but so far, that isn't helping.

OpenStudy (anonymous):

yea I got stuck there :(

OpenStudy (phi):

there must be a way

OpenStudy (anonymous):

using substitution method?

OpenStudy (mathmath333):

can we substitute 2theta=x

OpenStudy (anonymous):

why can't we just use the double angel formula and convert sin 2x into 2sinxcosx?

OpenStudy (phi):

how about \[ \frac{1}{2} \int \frac{\sin^2(x) + \cos^2(x)}{\sin(x) \cos(x)} dx \]

OpenStudy (phi):

that looks doable.

OpenStudy (phi):

write it as \[ \frac{1}{2} \int \frac{\sin(x)}{\cos(x)} + \frac{\cos(x)}{\sin(x)} \ dx \]

OpenStudy (anonymous):

YEAHHHH I got the answer already thank you very much you guys are so helpful glad I found you guys

OpenStudy (tkhunny):

Seriously, did you try \(u=\cot(2\theta)\)?